Given a number 𝑑 = 𝑑𝑒𝑑(𝐴1, β‹― , 𝐴𝑛 ), find 𝑑𝑒𝑑(𝐴𝑛, β‹― , 𝐴1 ).



1
Expert's answer
2021-08-03T16:32:09-0400

Property of determinant


"det(A_1, ..., A_i, ..., A_j, ..., A_n)=-det(A_1, ..., A_j, ..., A_i ..., A_n)"

We swap "A_1" and "A_n," "A_2" and "A_{n-1},..."

If "n=4k-3" or "n=4k,k=1,2,...," then


"det(A_1, ..., A_n)=det(A_n,...A_1)"

If "n=4k-2" or "n=4k-1,k=1,2,...," then


"det(A_1, ..., A_n)=-det(A_n,...A_1)"